OpenStudy (amberofficial1):

A. Use the following information to write an equation in slope-intercept form: 1. m = -2, b = 4 2. m = 1/2, b = -8 B. In the equation y = 1/4x + 9 3. What is the slope 4. What is the y-intercept 5. Explain how you would graph this equation

OpenStudy (flamingpiano):

Slope intercept is y=mx+b, so plug in the first #s and it gives you y=-2x+4, do the same for #2. M=slope, so use that for #3.

OpenStudy (flamingpiano):

b=y-intercept, for #4.

OpenStudy (amberofficial1):

that honestly makes no sense, im just trynna get answers here lol . i honestly dont know how to do this .

OpenStudy (flamingpiano):

Have you ever worked with slope before?

OpenStudy (anonymous):

This is the equation, y=mx+b All you have to do to figure out what the equation looks like is plug in the variables. So for number 1, the variables are m=-2 & b=4. So i just plug in those varibles into y=mx+b. It will turn out to be y=-2x+4

OpenStudy (flamingpiano):

Here's a more detailed answer. All of these questions are using the equation, slope-intercept, Y=mx+b. For #1, you plug in the -2 and 4 to get y=-2x+4. For #2, y=1/2x-8. Understand so far?

OpenStudy (anonymous):

The slope is m. The y-intercept is b.

OpenStudy (flamingpiano):

Yep,

OpenStudy (flamingpiano):

So for B, the slope would be 1/4, because that's in the position of m.